OpenStudy (anonymous):

Having difficulty with this question, please help. Perform the indicated operation and if possible, simplify, please show all of your work. Write your final answer as a fraction in lowest terms. 21/26 divided by 27/28.

OpenStudy (anonymous):

first step is to turn the division into a multiplication via \[\frac{21}{26}\div\frac{27}{28}=\frac{21}{26}\times \frac{28}{27}\]

OpenStudy (anonymous):

second step is to do any canceling you may find. for example you can cancel a 2 from 26 and 28, and a 3 from 21 and 27 to get \[\frac{7\times 18}{13\times 9}\] last step is to multiply

OpenStudy (anonymous):

28/2 is 14, not 18 :)

OpenStudy (anonymous):

so it would be (7x14)/(13x9) :)

OpenStudy (anonymous):

oops, so it is!!

OpenStudy (anonymous):

haha :)